OpenStudy (anonymous):

what is the value of 2q-8 when q = 6

OpenStudy (anonymous):

2(6)-8 12-8 4

OpenStudy (hoblos):

2q-8 put 6 in the place of q 2(6)-8 = 12-8=4

OpenStudy (anonymous):

2 x 6- 8 = 4